OpenStudy (anonymous):

-4(x + 1) - 3 = -3(x - 4)

OpenStudy (nurali):

-4(x + 1) - 3 = -3(x - 4) -4x-4-3=-3x+12 -4x-7=-3x+12 -4x+3x=12+7 -x=19 x=-19

OpenStudy (nurali):

Welcome to OpenStudy :D You may want to read the code of conduct ( http://openstudy.com/code-of-conduct ). Don't forget to click on the button called "Best Answer" to give a medal to the person who helped you :)

OpenStudy (mokeira):

Hahaha.... @Nurali you know the code and you knwo u shouldnt give answers just explanation